OpenStudy (anonymous):

Find the first three iterates of f(x)=x(3-x) if x0= -3 MULTIPLE CHOICE! A. x1=-18, x2=-378, x3=-144018 B. x1=0, x2=0, x3=0 C. x1=-18, x2=-270, x3=-73710 D. x1=18, x2=-270, x3=-73710

OpenStudy (anonymous):

My answer was C but its doesn't seem right

OpenStudy (phi):

first one is -3 * (3- (-3))= -3*(3+3)= -3*6 = -18 so choice A or C next one is -18* (3 - -18)= -18*(21)= - 378 that narrows it down to choice A. No need to figure out x3 (thank goodness)

OpenStudy (anonymous):

Thanks so much I see my mistake now!